OpenStudy (anonymous):

For the table below, determine the function rule, and find each of the missing y-values. Function Rule: ? x y 1 4 3 6 4 7 5 ? 7 10 10 ?

OpenStudy (anonymous):

well the answers i believe are 8 and 13 ...i notice what ever is added on one side is added on the other .... 1 +2 -= 3 4 +2 is 6 , 3 + 1 is 4 6+ 1 is 7 ..function rule i am not sure about though

OpenStudy (anonymous):

x=1,4,5,7,10 y=4,6,8,10,13 y=x+3

OpenStudy (anonymous):

y=x+3 ...

OpenStudy (anonymous):

that's the function rule?

OpenStudy (anonymous):

i don't think x + 3 is correct ..bc 6 = 4 +3 ..doesn't work

OpenStudy (anonymous):

ok.. so what's the rule?

OpenStudy (anonymous):

yes it is y=x+3 On my First post i messed up dont read that one actually im gonna delete it... But 1=4, 3=6, and 4=7. each of those all you do is add 3 and as you can see the numbers on the left are "x" and the numbers on the right are "y" , therefore y=x+3 ... get it now ?

OpenStudy (anonymous):

yep, thanks!

OpenStudy (anonymous):

Can i get a medel ? :D plz